# Semi-empirical constraints on the HI mass function of star-forming galaxies and $\Omega_{\rm HI}$ at $z\sim 0.37$ from interferometric surveys

### v1

@@ Appendix E The impact of sample size
Figure 9: Results for the HIMF from the combined stacking case, as a function of the chosen sample size $N$ and assuming the polynomial model for the $M_{\rm HI}-M_{\star}$ relation. We show the main results from this paper ($N=6,598$) as blue squares, for $N=10^{2}$ as gray circles, for $N=10^{3}$ as purple downward triangles, for $N=10^{4}$ as green stars, and for $N=10^{5}$ as orange diamonds. We also display the HIMF from ALFALFA at $z\sim 0$ as a black dashed line, for comparison. A horizontal tiny shift was applied to the data points to avoid overlaps and to ease visualization.
To assess the impact of the assumed sample size, we repeat the main procedure outline in Sect. 4, but varying the sample size. In this way, we aim at testing the impact on the HIMF of the sample size which we matched to the one used in the analysis from Bianchetti et al. (2025). Fig. 9 shows the results of this test. In particular, we compare the results for the combined stacking case from the ‘true’ sample size already presented in Sect. 5.2 (blue squares), to the resulting HIMF from increasingly larger samples of sizes $N=10^{2}$ (gray circles), $N=10^{3}$ (purple downward triangles), $N=10^{4}$ (green stars), and $N=10^{5}$ (orange diamonds). Ones sees that there are no evident systematic effects due to sample size for $N>1000$. Only in the case $N=100$ and $N=1000$ a small systematic shift of the HIMF for the $N=100$ is observed, although well within the uncertainties. In particular, for $N=100$ the HIMF is not measured for the used binning and setup at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\gtrsim 10.2$ due to an insufficient sampling.
@@ Appendix F Posterior distributions of the model parameters
Figure 10: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from assuming a linear model for the $M_{\rm HI}-M_{\star}$ relation
Figure 11: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from assuming a polynomial model for the $M_{\rm HI}-M_{\star}$ relation
Figure 12: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from the combined stacking case. A polynomial model for the $M_{\rm HI}-M_{\star}$ relation was assumed.
Figure 13: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from the MIGHTEE stacking case. A polynomial model for the $M_{\rm HI}-M_{\star}$ relation was assumed.
Figure 14: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from the CHILES stacking case. A polynomial model for the $M_{\rm HI}-M_{\star}$ relation was assumed.

### v2

@@ Appendix E The impact of sample size
To assess the impact of the assumed sample size, we repeat the main procedure outline in Sect. 4, but varying the sample size. In this way, we aim at testing the impact on the HIMF of the sample size which we matched to the one used in the analysis from Bianchetti et al. (2025). Fig. 9 shows the results of this test. In particular, we compare the results for the combined stacking case from the ‘true’ sample size already presented in Sect. 5.2 (blue squares), to the resulting HIMF from increasingly larger samples of sizes $N=10^{2}$ (gray circles), $N=10^{3}$ (purple downward triangles), $N=10^{4}$ (green stars), and $N=10^{5}$ (orange diamonds). One sees that there are no evident systematic effects due to sample size for $N>1000$. Only in the case $N=100$ and $N=1000$ a small systematic shift of the HIMF is observed, although well within the uncertainties. In particular, for $N=100$ the HIMF is not measured for the used binning and setup at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\gtrsim 10.2$ due to an insufficient sampling.
@@
Figure 9: Results for the HIMF from the combined stacking case, as a function of the chosen sample size NN and assuming the polynomial model for the $M_{\rm HI}-M_{\star}$ relation. We show the main results from this paper ($N=6598$) as blue squares, for $N=10^{2}$ as gray circles, for $N=10^{3}$ as purple downward triangles, for $N=10^{4}$ as green stars, and for $N=10^{5}$ as orange diamonds. We also display the HIMF from ALFALFA at $z\sim 0$ as a black dashed line, for comparison. A horizontal tiny shift was applied to the data points to avoid overlaps and to ease visualization.
@@ Appendix F Posterior distributions of the model parameters
Figure 10: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from assuming a linear model for the $M_{\rm HI}-M_{\star}$ relation. The contours represent the $68\%$ and $99\%$ confidence levels.
@@
Figure 11: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from assuming a polynomial model for the $M_{\rm HI}-M_{\star}$ relation. The contours represent the $68\%$ and $99\%$ confidence levels.
Figure 12: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from the combined stacking case. The contours represent the $68\%$ and $99\%$ confidence levels.
Figure 13: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from the MIGHTEE stacking case. The contours represent the $68\%$ and $99\%$ confidence levels.
Figure 14: Posterior distributions of the best-fitting parameters of the Schechter profile to the HIMF, resulting from the CHILES stacking case. The contours represent the $68\%$ and $99\%$ confidence levels.